OpenStudy (anonymous):

a 10-foot ladder is leaning against a wal. If a person pulls the base of the ladder away from the wall at the rate of 1 foot per second, how fast is the top going down when the base of the ladder is 6 feet from the wall?

OpenStudy (anonymous):

@ganeshie8 @hartnn The roc in movement is the derivative, right?

OpenStudy (anonymous):

@mathmale Is this a good way to look at it?

OpenStudy (anonymous):

I've drawn a triangle to represent it

OpenStudy (anonymous):



OpenStudy (anonymous):

and x should equal 6, right?

OpenStudy (anonymous):

we are given the change of x with respect to time which we can write as dx/dt = 6

OpenStudy (anonymous):

nope, they are giving you a rate

OpenStudy (anonymous):

Ah yes, of course, because it is changing with respect to time

OpenStudy (anonymous):

so think hmmm change of distance over change of time.... if distance is x and time is t then we write dx/dt

OpenStudy (anonymous):

6 is the final pos it is approaching

OpenStudy (anonymous):

yes that is usually the tricky part... putting their words into derivative form

OpenStudy (anonymous):

oh gosh I am sorry I must be too tired

OpenStudy (anonymous):

you are correct, dx/dt is 1ft/s

OpenStudy (anonymous):

hmm, ok so we have some intervals, but I'm not sure how to use this info

OpenStudy (anonymous):

okay well we have a right angle triangle... what do we know that relates all the sides

OpenStudy (anonymous):

now when the ladder moves away at 1ft/s, the top should move at 1ft/s too then?

OpenStudy (anonymous):

x^2+y^2=10^2

OpenStudy (anonymous):

hmm wait yes I was about to say use pyth.

OpenStudy (anonymous):

and just to make sure you know about this... the reason I am subbing in the 10feet now is because it is the length of the ladder, which is a constant (the ladder isn't getting longer or shorter)

OpenStudy (anonymous):

yes, it makes sense, the others vary over time

OpenStudy (anonymous):

Okay so step 1. find all variables 2. find an equation that relates all the variables (phythagorus works here!) 3. find the derivative of that equation with respect to time

OpenStudy (anonymous):

because we already know dx/dt... so if we take the derivative of respect to time we can just sub that in!

OpenStudy (anonymous):

do I need to calculate the distance from the top of the ladder to the ground for each interval?

OpenStudy (anonymous):

hmm I am not quite sure

OpenStudy (anonymous):

nope they just want to know when the base of the ladder is 6 feet from the wall (so when x=6)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do lets take the derivative of x^2+y^2=0 with respect to time

OpenStudy (anonymous):

that should be a 10^2 at the end there

OpenStudy (anonymous):

wait why 0?

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

heh actually it wouldn't matter because the derivative is 0 anyways

OpenStudy (anonymous):

yes, is it 2x/dt +2y/dt =0 ?

OpenStudy (anonymous):

Just a bit more info about that 10 thing (the length of the ladder). If you forgot to sub in the 10 and just made it a variable (z) you would find the derivative to be (2z)(dz/dt)

OpenStudy (anonymous):

hmm almost... is should be (2x)(dx/dt) + (2y)(dy/dt) = 0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but yeah, what I was saying about the (2z)(dz/dt). If you end up doing this and start to panic because you don't know either, just think that the ladder (z) is not changing, so (dz/dt)=0 and so (2z)(dz/dt)=0

OpenStudy (anonymous):

ok, Ill remember that thanks

OpenStudy (anonymous):

I hope this isn't confusing. This really messed me up when I started doing these questions... I was like when do I plug in the values??? now or later???

OpenStudy (anonymous):

)

OpenStudy (anonymous):

so did you see where you went wrong? when you take the derivative with respect to time of x^2 you must use the chain rule (2x)(dx/dt)

OpenStudy (anonymous):

hmm because it is x*x

OpenStudy (anonymous):

2 chain rules and a sum rule

OpenStudy (anonymous):

I guess you could look at it like that

OpenStudy (anonymous):

I think of it as start with the outside (the 2 exponent) and move it in front of the x. Then you must take the derivative of the x with respect to t thus 2x(dx/dt)

OpenStudy (anonymous):

this is confusing at the start because up till you learn implicit differentiation you were taking the derivative with respect to x. so (d/dx) of (x^2) is 2x(dx/dx) and the dx cancels out leaving 2x

OpenStudy (anonymous):

ah, the way I did it was: g=x g'=1 f=u^2 f'=2u

OpenStudy (anonymous):

yes we need to keep in mind we are taking the derivative with respect to time

OpenStudy (anonymous):

so do you understand how to get to 2x(dx/dt) + 2y(dy/dt) = 0 now?

OpenStudy (anonymous):

then used the chain rule: 2(x)(1) and then the dx/dt confused me

OpenStudy (anonymous):

let me look over what you said

OpenStudy (anonymous):

why did you put that 1 there

OpenStudy (anonymous):

with the 2(x)(1)

OpenStudy (anonymous):

derivative of x is 1

OpenStudy (anonymous):

g=x g'=1 f=u^2 f'=2u

OpenStudy (anonymous):

the derivative of x with RESPECT TO x might be 1.... but we want the derivative of x with RESPECT TO t

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

this factor still confuses me

OpenStudy (anonymous):

up until now you have probably been asked to find dy/dx=blah blah blah right

OpenStudy (anonymous):

more or less

OpenStudy (anonymous):

I used to hate that notation and always used y' or whatever

OpenStudy (anonymous):

but the dy/dx shows you what you are taking the derivative of with respect to

OpenStudy (anonymous):

using the chain rule: (d/dt) (x^2) first move the 2 over to the front of the x THEN you must multiply that by the DERIVATIVE OF THE INSIDE FUNCTION WITH RESPECT TO TIME!

OpenStudy (anonymous):

I dislike those words " with respect to" >_<

OpenStudy (anonymous):

hehe it just means what is x doing when t is whatever

OpenStudy (anonymous):

when you take dy/dx, you are finding out what y is doing when x is doing something else... there is probably a better way to explain this

OpenStudy (anonymous):

so the inside function is x

OpenStudy (anonymous):

so x times d/dt is dx/dt

OpenStudy (anonymous):

yes... and up till now you have been finding the derivative of x with respect to x (which means what is x doing when x is doing something... well they are the same variable! as x increase so does x so it would just be in a 1/1 ratio simplifying to 1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which explains why it's not always written

OpenStudy (anonymous):

don't just think of it as times though... think of it as what is this variable (x) doing when t is doing something else

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes because at the beginning they just want you to get familiar with all the rules

OpenStudy (anonymous):

I guess I mean what you are doing arithmetically to get this result using the chain rule

OpenStudy (anonymous):

What sometimes helps me is to try using these rules on simple problems that I already know I can solve other ways... it proves to me that these other weirder ways work

OpenStudy (anonymous):

yep, just the chain rule

OpenStudy (anonymous):

and when you get to a variable that you don't know the answer to, just put it with respect to the other variable and move on!

OpenStudy (anonymous):

but this dx/dt is our g' in the chain rule then, right?

OpenStudy (anonymous):

so back to x^2+y^2=10^2

OpenStudy (anonymous):

what is the formula for chain rule again

OpenStudy (anonymous):

just to clarify

OpenStudy (anonymous):

ah yes d/dx(fog)(x)=d/dx f(g(x))=f'(g(x)) g'(x)

OpenStudy (anonymous):

I would say that g'(x) = dx/dt in our situation

OpenStudy (anonymous):

ok :) cool, that's how I visualized it, ok, I understand this part, thanks

OpenStudy (anonymous):

good so we have 2x(dx/dt) + 2y(dy/dt) = 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can sub in dx/dt, right?

OpenStudy (anonymous):

lets just remember what each thing is x is the length the ladder is away from the wall dx/dt is the rate that the ladder is moving away from the wall y is the height of the ladder from the ground dy/dt is the rate that the ladder is moving down the wall